how many grams are contained in 1.52 X 10^36 particles of MgO


1
Expert's answer
2021-02-05T04:02:41-0500

According to Avogadro's number, 1 mol of any substance contains 6.022 × 1023 particles. From here:

n (MgO) = 1.52 × 1036 × 1 mol / 6.022 ×1023 = 2.5 × 1012 mol

According to the equation:

n = m / Mr

where n - number of moles, m - mass, Mr - molecular weight

From here:

m = n × Mr = 2.5 × 1012 mol × 40 g/mol = 1014 g


Answer: 1014 g